OpenStudy (anonymous):

in a family with 3 children excluding multiple births what is the probability of having 3 girls? Assume that a girl is as likely as a boy at each birth.

OpenStudy (anonymous):

My random guess of 1/3. LOL.

OpenStudy (anonymous):

probability of having a girl, so you are told, is \(\frac{1}{2}\) so the probability of having 3 in a row is \(\left(\frac{1}{2}\right)^3\)

OpenStudy (blockcolder):

What he said.